# The large sieve for square moduli, revisited

### v3

@@ 9.2. Treatment of the general case
Since $V_{h}$ has compact support in $[C_{0},C_{1}]\subset\mathbb{R}_{>0}$ and $h/(Q^{2}rz)\ll 1$, we have
@@ 10.1. Sketch for our particular situation
An analogous treatment of general $Q$, $N$ and $z$ in the following subsection will lead to a non-trivial bound for $P(x)$ in an $r$-range which in the case of (39) takes the form $r\leq Q^{1/4-\eta}$ for some $\eta>0$.
@@ 12. An interesting connection to a short character sum
where $\nu:=R/r$ and $\Phi$ is a Schwartz class function satisfying $0\leq\Phi\leq 1$ and supported in $[-1,1]$. We define the corresponding weighted additive energy as
@@
$E_{2}^{\prime}(R;j,r)=\sum\limits_{\begin{subarray}{c}(k_{1},k_{2},k_{3},k_{4})\bmod r\\ k_{1}+k_{2}\equiv k_{3}+k_{4}\bmod{r}\\ \left|\left|k_{i}^{2}/r\right|\right|\leq R/r\text{ for }i=1,2,3,4\end{subarray}}1=\sum\limits_{\begin{subarray}{c}(k_{1},k_{2},k_{3},k_{4})\bmod r\\ k_{1}+k_{2}\equiv k_{3}+k_{4}\bmod{r}\end{subarray}}\prod\limits_{i=1}^{4}\chi_{[0,R/r]}\left(\left|\left|\frac{\overline{j}k_{i}^{2}}{r}\right|\right|\right).$
The only differences between the definitions of $E_{2}(R;j,r)$ and $E_{2}^{\prime}(R;j,r)$ are that in the definition of $E_{2}^{\prime}(R;j,r)$, the fractional part is replaced by the distance to the nearest integer, and $k_{i}$’s congruent to 00 modulo $r$ are no longer excluded. In the case of prime moduli $r$, the method of Kerr, Shkredov, Shparlinski and Zaharescu in [22] gives the same bound for $E_{2}^{\prime}(R;j,r)$ as the bound (10) for $E_{2}(R;j,r)$, namely
@@
where the last inequality arises from the rapid decay of $\Phi$. Applying Poisson summation, Proposition 10 with the modulus 1, $L=\nu$ and $M=-y$, we have
@@
$\delta(a_{1},a_{2},l):=\begin{cases}1&\mbox{ if }a_{1}\equiv a_{2}\equiv 0\bmod{r}\mbox{ and }l\not\equiv 0\bmod{r},\\ 1&\mbox{ if }a_{1}\equiv-a_{2}\bmod{r}\mbox{ and }l\equiv 0\bmod{r},\\ 0&\mbox{ otherwise.}\end{cases}$
@@
If $r$ is an odd prime, $1\leq M\leq r$ and $W:\mathbb{R}\rightarrow\mathbb{C}$ is a Schwartz class function whose Fourier transform satisfies $0\leq\hat{W}\leq 1$ and is supported in $[-1,1]$, then
@@
If $r$ is an odd prime and $W:\mathbb{R}\rightarrow\mathbb{C}$ is a Schwartz class function whose Fourier transform satisfies $0\leq\hat{W}\leq 1$ and is supported in $[-1,1]$, then
@@ 14. Appendix: Estimation of a complete exponential sum
Since $(2js,p)=1$, we have $\mbox{ord}_{p}(f^{\prime})=t$, where $p^{t}=(bu^{2},a,bk^{2},p^{m})$. If $t=m$, then (130) holds trivially. If $t=m-1$, then dividing $f(x)$ by $p^{t}$, we obtain an exponential sum modulo $p$ which can be estimated by Proposition 13, thus establishing (130) as well. If $p\not=2$ and $m\geq t+2$, then Proposition 14 implies (130) since the number of roots of $p^{-t}f^{\prime}(x)$ modulo $p$ is bounded by $4$ and their multiplicities as well. Using (128) and (130), we deduce that

### v4

@@ 9.2. Treatment of the general case
Since $V_{h}$ has compact support in $C_{0},C_{1}]\subset\mathbb{R}_{>0}$ and $h/(Q^{2}rz)\ll 1$, we have
@@ 10.1. Sketch for our particular situation
An analogous treatment of general $Q$, NN and $z$ in the following subsection will lead to a non-trivial bound for $P(x)$ in an $r$-range which in the case of (39) takes the form $r\leq Q^{1/4-\eta}$ for some $\eta>0$.
@@ 12. An interesting connection to a short character sum
where $\nu:=R/r$ and $\Phi$ is a Schwartz class function satisfying $0\leq\Phi\leq 1$ and supported in [1,1][-1,1]. We define the corresponding weighted additive energy as
@@
[…]
The only differences between the definitions of $E_{2}(R;j,r)$ and $E_{2}^{\prime}(R;j,r)$ are that in the definition of $E_{2}^{\prime}(R;j,r)$, the fractional part is replaced by the distance to the nearest integer, and $k_{i}$’s congruent to 0 modulo $r$ are no longer excluded. In the case of prime moduli $r$, the method of Kerr, Shkredov, Shparlinski and Zaharescu in [23] gives the same bound for $E_{2}^{\prime}(R;j,r)$ as the bound (10) for $E_{2}(R;j,r)$, namely
@@
where the last inequality arises from the rapid decay of $\Phi$. Applying Poisson summation, Proposition 10 with the modulus 1, $L=\nu$ and M=yM=-y, we have
@@
[…]
@@
If $r$ is an odd prime, $1\leq M\leq r$ and $W:\mathbb{R}\rightarrow\mathbb{C}$ is a Schwartz class function whose Fourier transform satisfies $0\leq\hat{W}\leq 1$ and is supported in [1,1][-1,1], then
@@
If $r$ is an odd prime and $W:\mathbb{R}\rightarrow\mathbb{C}$ is a Schwartz class function whose Fourier transform satisfies $0\leq\hat{W}\leq 1$ and is supported in [1,1][-1,1], then
@@ 14. Appendix: Estimation of a complete exponential sum
Since $(2js,p)=1$, we have $\mbox{ord}_{p}(f^{\prime})=t$, where $p^{t}=(bu^{2},a,bk^{2},p^{m})$. If $t=m$, then (130) holds trivially. If t=m1t=m-1, then dividing $f(x)$ by $p^{t}$, we obtain an exponential sum modulo $p$ which can be estimated by Proposition 13, thus establishing (130) as well. If $p\not=2$ and $m\geq t+2$, then Proposition 14 implies (130) since the number of roots of $p^{-t}f^{\prime}(x)$ modulo $p$ is bounded by $4$ and their multiplicities as well. Using (128) and (130), we deduce that